# Monte Carlo Simulation of Operator Dynamics and Entanglement in Dual-Unitary Circuits

### v3

@@ III.2.3 Case 3: from the left light-cone edge, focusing on left-movers only
We can construct a scenario where an entanglement transition can be observed at a finite $\alpha$ within the physical range of the gate parameter. This is described in Case 3, where we consider only the subsystem $A$ of left movers. Following the analysis of Case 2, the only difference is a reduction of $\overline{S^{(n)}_{\tilde{A},1}}$ by a factor of two. Therefore, we expect
$e^{-(n-1)\overline{S^{(n)}_{A,3}}}\approx\alpha^{n}\sum_{1\leq t_{0}\leq l_{A}+1}(1-\alpha)^{n(t_{0}-1)}4^{-(n-1)(l_{A}-t_{0}+1)}.$ (38)
This results in a change in the critical $\alpha$:
@@
III.2.4 Summary of results on operator entanglement
In this subsection, we studied the dynamics of operator entanglement in dual-unitary circuits, examining three distinct cases based on different subregions near the light-cone with the initial operator $O=Z_{1}$, as illustrated in Fig. 8. Across all three cases, the $n$th-Rényi operator entanglement entropy $\overline{S_{A}^{(n)}}(t)$, as defined in Eq. (13), exhibits a volume-law scaling on both edges of the light cone in the long-time limit $t\to\infty$, i.e.
$\overline{S_{A}^{(n)}}\simeq a^{(n)}l_{A}+\cdots,$ (40)
growing linearly with the subregion size $l_{A}$. However, the dependence of the volume-law coefficient $a^{(n)}$ on the gate parameter $\alpha$ and the Rényi index $n$ differs between the cases, as summarized in Fig. 12.
@@ III.2.4 Summary of results on operator entanglement
The transition from maximal to sub-maximal volume-law growth in operator entanglement is observed specifically in Case 3, when focusing on the left movers in the left light-cone entanglement region. This result highlights how the gate parameter $\alpha$ influence the degree of operator entanglement among the leading left movers emitted from the expanding operator. Such operator entanglement transitions, characterized by the abrupt drop in the volume-law coefficient at a critical value of $\alpha$, have been previously discussed in the literature [40, 37, 36]. Notably, Ref. [37] first identified the operator entanglement transition in a setup akin to our Case 2 and derived the following formula for qubit systems:
[…]
where $\lambda$ is the leading non-trivial eigenvalue of a transfer matrix defined therein. However, the explicit dependence of $\lambda$ on the gate parameter was not fully established. Our study builds upon this foundation by demonstrating that this behavior actually aligns precisely with our Case 3 scenario when we identify $|\lambda|^{2}=1-\alpha=1-\frac{2}{3}\cos^{2}(2J)$.
@@ IV Discussion
Acknowledgements.
Acknowledgments —  We thank Yimu Bao, Tian-Gang Zhou for discussions. MHS, TTW and ZYM acknowledge the support from the Research Grants Council (RGC) of Hong Kong (Project Nos. 17301721, AoE/P-701/20, 17309822, HKU C7037-22GF, 17302223, 17301924), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A_HKU703/22), the GD-NSF (No. 2022A1515011007). PZ is supported by the Shanghai Rising-Star Program under grant number 24QA2700300, the NSFC under grant 12374477, and the Innovation Program for Quantum Science and Technology ZD0220240101. YZY is supported by a startup fund from UCSD. We thank HPC2021 system under the Information Technology Services, The University of Hong Kong, as well as the Beijing PARATERA Tech CO.,Ltd. (URL: https://cloud.paratera.com) for providing HPC resources that have contributed to the research results reported within this paper. ZYZ, YZY and ZYM acknowledge the hospitality of Kavli Institute for Theoretical Physics (KITP) and thank the organizers of the KITP program “Correlated Gapless Quantum Matter” where this work was initiated. This research was supported in part by grant NSF PHY-2309135 to the KITP.

### v4

@@ 3.2.3 Case 3: from the left light-cone edge, focusing on left-movers only
We can construct a scenario where an entanglement transition can be observed at a finite $\alpha$ within the physical range of the gate parameter. This is described in Case 3, where we consider only the subsystem $A$ of left movers. Following the analysis of Case 2, the only difference is a reduction of $\overline{S^{(n)}_{\tilde{A},1}}$ by a factor of two. Therefore, we expect {myequation} e^-(n-1)¯S^(n)_A,3≈α^n∑_1≤t_0≤l_A+1(1-α)^n(t_0-1)4^-(n-1)(l_A-t_0+1). This results in a change in the critical $\alpha$:
@@
3.2.4 Summary of results on operator entanglement
In this subsection, we studied the dynamics of operator entanglement in dual-unitary circuits, examining three distinct cases based on different subregions near the light-cone with the initial operator $O=Z_{1}$, as illustrated in Fig. 8. Across all three cases, the $n$th-Rényi operator entanglement entropy $\overline{S_{A}^{(n)}}(t)$, as defined in Eq. (11), exhibits a volume-law scaling on both edges of the light cone in the long-time limit $t\to\infty$, i.e. {myequation} ¯S_A^(n)≃a^(n)l_A+⋯, growing linearly with the subregion size $l_{A}$. However, the dependence of the volume-law coefficient $a^{(n)}$ on the gate parameter $\alpha$ and the Rényi index $n$ differs between the cases, as summarized in Fig. 12.
@@ 3.2.4 Summary of results on operator entanglement
The transition from maximal to sub-maximal volume-law growth in operator entanglement is observed specifically in Case 3, when focusing on the left movers in the left light-cone entanglement region. This result highlights how the gate parameter $\alpha$ influence the degree of operator entanglement among the leading left movers emitted from the expanding operator. Such operator entanglement transitions, characterized by the abrupt drop in the volume-law coefficient at a critical value of $\alpha$, have been previously discussed in the literature [40, 37, 36]. Notably, Ref. [37] first identified the operator entanglement transition in a setup akin to our Case 2 and derived the following formula for qubit systems: {myequation} a^(n)= {-nn-1ln|λ|2if $1>|\lambda|^{2}>2^{-\frac{2(n-1)}{n}}$,2ln2otherwise, where $\lambda$ is the leading non-trivial eigenvalue of a transfer matrix defined therein. However, the explicit dependence of $\lambda$ on the gate parameter was not fully established. Our study builds upon this foundation by demonstrating that this behavior actually aligns precisely with our Case 3 scenario when we identify $|\lambda|^{2}=1-\alpha=1-\frac{2}{3}\cos^{2}(2J)$.
@@ 4 Discussion
We thank Yimu Bao, Tian-Gang Zhou for discussions. MHS, TTW and ZYM acknowledge the support from the Research Grants Council (RGC) of Hong Kong (Project Nos. 17301721, AoE/P-701/20, 17309822, HKU C7037-22GF, 17302223, 17301924), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A_HKU703/22), the GD-NSF (No. 2022A1515011007). PZ is supported by the Shanghai Rising-Star Program under grant number 24QA2700300, the NSFC under grant 12374477, and the Innovation Program for Quantum Science and Technology 2024ZD0300101. YZY is supported by a startup fund from UCSD. We thank HPC2021 system under the Information Technology Services, The University of Hong Kong, as well as the Beijing PARATERA Tech CO.,Ltd. (URL: https://cloud.paratera.com) for providing HPC resources that have contributed to the research results reported within this paper. ZYZ, YZY and ZYM acknowledge the hospitality of Kavli Institute for Theoretical Physics (KITP) and thank the organizers of the KITP program “Correlated Gapless Quantum Matter” where this work was initiated. This research was supported in part by grant NSF PHY-2309135 to the KITP.